OpenStudy (anonymous):

What is the sum of a 12-term arithmetic sequence where the last term is 13 and the common difference is -10

OpenStudy (anonymous):

810

OpenStudy (anonymous):

are you sure its not 816?

OpenStudy (anonymous):

a12=a1+11d 13=a1+(-110) a1=123 S=6(123+13) =816

OpenStudy (anonymous):

oky thats what thought, thank you !

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

oh sorry.i forgot to add 6 at the last step of my calculation!!